OpenStudy (anonymous):

Verify the identity. cos 4x + cos 2x = 2 - 2 sin2 2x - 2 sin2 x

OpenStudy (loser66):

I don't get why "verify" and then "increasing" "decreasing" hey, why do you delete your comments?

OpenStudy (anonymous):

because it wasnt related to the question

OpenStudy (ybarrap):

\(\text{Proof that }\cos4x+\cos2x=2-2\sin^2 2x-2\sin^2 x\\\) \(=2-2\left [{1-\cos4x\over2} \right ]-2\left [{1-\cos2x\over2} \right ]\\\) \(=2-\left (1-\cos4x \right )-2\left (1-\cos2x \right )\\\) \(=\cos4x +\cos2x\\\)

OpenStudy (ybarrap):

Let me know if you have any questions.

OpenStudy (anonymous):

did you use a formula

OpenStudy (ybarrap):

The only formula used was the half-angle formula: \(\large \sin^2 x={1-\cos2x\over2}\). Have you seen this?

OpenStudy (anonymous):

no i havent